OpenStudy (anonymous):

Explain how to graph x – 2y = 4 using the x- and y-intercepts method. Be sure to use complete sentences and use the x- and y-intercepts method only.

OpenStudy (anonymous):

First find the x and y intercepts. To find the y intercept, plug in 0 for x and solve for y. Then graph the point (0,y) (y is the value you just solved for, Then find the x intercept. Plug in 0 for y and solve for x. Graph the point (x,0) x is the value you just solved for. you now have 2 points to make a line.

OpenStudy (anonymous):

-2y=4-x y=-2+x draw graph, with slope of 1 and you will get the values for X and Y